# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 4.1. Lipschitz estimates for homogeneous systems
$|D_{x^{\prime}}u|+\sqrt{\lambda}|u|\leq NC_{2}.$
To estimate $D_{d}u$, note that $|D_{d}u|\leq N(|\mathcal{U}|+|D_{x^{\prime}}u|)$. From the definition of $\mathcal{U}$ and the equation, we have $|\partial_{t}^{I}D_{x^{\prime}}^{J}\mathcal{U}|\leq N|D\partial_{t}^{I}D_{x^{\prime}}^{J}u|$ and $|D_{d}\partial_{t}^{I}D_{x^{\prime}}^{J}\mathcal{U}|\leq N(|\partial_{t}\partial_{t}^{I}D_{x^{\prime}}^{J}u|+|DD_{x^{\prime}}\partial_{t}^{I}D_{x^{\prime}}^{J}u|+\lambda|\partial_{t}^{I}D_{x^{\prime}}^{J}u|)$ for $I,J\geq 0$. We then estimate the term $\mathcal{U}$ similarly as in the estimate of $u$ above, and get
$|\mathcal{U}|\leq NC_{2}.$
Combining the above estimates, we arrive to (4.6).
By using an iteration argument as in [25, p. 184-186], estimates (4.6) hold for general $p\in(1,2]$, assuming that initially $u\in\mathcal{H}^{1}_{2}(Q_{r})$. Assume now that $u\in\mathcal{H}^{1}_{p}(Q_{1})$, where $p\in(1,2)$. Let $u^{(\varepsilon)}$ be the mollification of $u$ in the $t,x^{\prime}$ variables. As above,
$\|u^{(\varepsilon)}\|_{L_{\infty}(Q_{3/4})}\lesssim\int_{-4/5}^{4/5}(\|D_{d}u^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q^{\prime}_{4/5})}+\|u^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q^{\prime}_{4/5})})ds<\infty$
for large enough even $k=k(p,d)$. Then $\|D_{x^{\prime}}u^{(\varepsilon)}\|_{L_{\infty}(Q_{3/4})}<\infty$ as well. Next,
$\|\mathcal{U}^{(\varepsilon)}\|_{L_{\infty}(Q_{3/4})}$ $\lesssim\int_{-4/5}^{4/5}(\|D_{d}\mathcal{U}^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q^{\prime}_{4/5})}+\|\mathcal{U}^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q^{\prime}_{4/5})})ds$
$\lesssim\int_{-4/5}^{4/5}(\|D_{d}u^{(\varepsilon)}(\cdot,s)\|_{W^{k/2+1,k+2}_{p}(Q^{\prime}_{4/5})}+\|u^{(\varepsilon)}(\cdot,s)\|_{W^{k/2+1,k+2}_{p}(Q^{\prime}_{4/5})})ds<\infty,$
and therefore $\|D_{d}u^{(\varepsilon)}\|_{L_{\infty}(Q_{3/4})}<\infty$ as well. In particular, $u^{(\varepsilon)}\in\mathcal{H}^{1}_{2}(Q_{3/4})$ for all small enough $\varepsilon$. Then, (4.6) holds with $u^{(\varepsilon)}$ in place of $u$, and with $3r/4$ in place of $r$ in the right hand sides. Passing to the limit as $\varepsilon\to 0^{+}$, we obtain (4.6) for $u$. ∎
In the following lemma, we prove boundary $L_{\infty}$ estimates of solutions.
@@
Let $p\in(1,2]$, $r>0$, and $u\in\mathcal{H}^{1}_{p}(Q_{r}^{+},\mu)$ be a weak solution to (4.4)-(4.5) in $Q_{r}^{+}$. Then
$\begin{gathered}\|u\|_{L_{\infty}(Q_{r/2}^{+})}\leq N\left(\fint_{Q_{r}^{+}}|u(t,x)|^{p}d\mu(z)\right)^{1/p},\\ \|Du\|_{L_{\infty}(Q_{r/2}^{+})}+\sqrt{\lambda}\|u\|_{L_{\infty}(Q_{r/2}^{+})}\leq N\left(\fint_{Q_{r}^{+}}(|Du(t,x)|^{p}+\lambda^{p/2}|u(t,x)|^{p})d\mu(z)\right)^{1/p},\\ \|\mathcal{U}/x_{d}\|_{L_{\infty}(Q_{r/2}^{+})}\leq Nr^{-1}\left(\fint_{Q_{r}^{+}}(|Du(t,x)|^{p}+\lambda^{p/2}|u(t,x)|^{p})d\mu(z)\right)^{1/p},\end{gathered}$ (4.7)
where $N=N(d,n,\alpha,p,\kappa)>0$.
@@
By scaling, it is sufficient to consider the case when $r=1$. We first consider the case when $p=2$. For compactness of notation, let
@@
Fix an even integer $k>(d+1)/2$. From the equation, we have $|D_{d}(x_{d}^{\alpha}\mathcal{U})|\leq Nx_{d}^{\alpha}(|u_{t}|+|DD_{x^{\prime}}u|+\lambda|u|)$. Then from $\lim_{x_{d}\to 0^{+}}x_{d}^{\alpha}\mathcal{U}=0$, the Sobolev embedding, Hölder’s inequality, and Lemma 4.3, for $z\in Q_{1/2}^{+}$ it holds that
@@
$\leq NC_{1}+N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})\left(\int_{0}^{1/2}\|D_{d}u(\cdot,s)\|^{2}_{W_{2}^{k/2,k}(Q_{1/2}^{\prime})}s^{\alpha}ds\right)^{1/2}\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})C_{1}.$
In the above, estimate $|u(z^{\prime},1/2)|\leq NC_{1}$ follows from the interior $L_{\infty}$ estimate in Lemma 4.4, $\varepsilon\in(0,1)$ is any constant when $\alpha\in\{1,3,5,\dots\}$, and $\varepsilon=0$ otherwise. We introduce small $\varepsilon>0$ to avoid the log-correction appearing in integrals $\int x_{d}^{-1}dx_{d}$. Applying estimate (4.9) to $u_{t},D_{x^{\prime}}u,D^{2}_{x^{\prime}}u,\lambda u$, and $\lambda^{1/2}u$, and using Lemma 4.3 give that on $Q_{1/2}^{+}$ we have
$\begin{gathered}|u_{t}|+|D_{x^{\prime}}u|+|D^{2}_{x^{\prime}}u|+\lambda|u|+\lambda^{1/2}|u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})C_{2}.\end{gathered}$ (4.10)
@@
$|D_{d}u|\leq N|\mathcal{U}|+N|D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})C_{2}.$ (4.11)
@@
$|D_{d}D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})C_{2}.$ (4.12)
@@
$\begin{gathered}|x_{d}^{\alpha}\mathcal{U}|\leq NC_{2}\int_{0}^{x_{d}}s^{\alpha}(1+s^{\frac{1-\alpha}{2}-\varepsilon})ds\leq N(x_{d}^{1+\alpha}+x_{d}^{\frac{3+\alpha}{2}-\varepsilon})C_{2},\end{gathered}$
@@
$|\mathcal{U}/x_{d}|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\varepsilon})C_{2}.$
If $\alpha\in(-1,1)$, we are done with the case $p=2$. If $\alpha\geq 1$, we iterate the estimates above to improve the exponent $\frac{1-\alpha}{2}-\varepsilon$ as follows. Plugging (4.11) into the first line of (4.9), we get
$|u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\varepsilon})C_{1}.$
@@
$|u_{t}|+|D_{x^{\prime}}u|+|D^{2}_{x^{\prime}}u|+\lambda|u|+\lambda^{1/2}|u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\varepsilon})C_{2}.$
@@
$|D_{d}u|+|D_{d}D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\varepsilon})C_{2}.$
@@
$|\mathcal{U}/x_{d}|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\varepsilon})C_{2}.$
@@
By using an iteration argument as in [25, p. 184-186], estimates (4.7) hold for general $p\in(1,2]$, assuming that initially $u\in\mathcal{H}^{1}_{2}(Q_{r}^{+},\mu)$. Assume now that $u\in\mathcal{H}^{1}_{p}(Q_{1}^{+},\mu)$, where $p\in(1,2)$. Let $u^{(\varepsilon)}$ be the mollification of $u$ in the $t,x^{\prime}$ variables. Note that the argument above can essentially be repeated by taking
$C_{1}=C_{2}=\left(\int_{0}^{4/5}\Bigl[\|Du^{(\varepsilon)}\|_{W^{k/2,k}_{p}(Q_{4/5}^{\prime})}^{p}+\|u^{(\varepsilon)}\|_{W^{k/2,k}_{p}(Q_{4/5}^{\prime})}^{p}\Bigr]s^{\alpha}ds\right)^{1/p}<\infty,$
where $k=k(p,d)$ is a large enough even number. Indeed, by following the derivation of estimates above, the following sequence of estimates can be obtained for $z\in Q_{3/4}^{+}$. Let $\delta\in(0,1)$ be any constant if $\frac{1+\alpha}{p}\in\{1,2,3,\dots\}$ and $\delta=0$ otherwise. Then
@@
and so on. In finitely many steps, we get $|u^{(\varepsilon)}|+|Du^{(\varepsilon)}|+|\mathcal{U}^{(\varepsilon)}/x_{d}|\in L_{\infty}(Q_{3/4}^{+})$. In particular, $u^{(\varepsilon)}\in\mathcal{H}^{1}_{2}(Q_{3/4}^{+},\mu)$ for all small enough $\varepsilon$. Then, (4.7) holds with $u^{(\varepsilon)}$ in place of $u$, and with $3r/4$ instead of $r$ in the right hand sides. Passing to the limit as $\varepsilon\to 0^{+}$, we obtain (4.7) for $u$. ∎
@@
It should be noted that in the scalar case $n=1$ estimates (4.7) were proved in [17, pp. 1481–1487] with a somewhat different approach. There the authors first used Moser’s iteration to show the bound $\|u\|_{L_{\infty}(Q_{1/2}^{+})}\leq N\|u\|_{L_{2}(Q_{1}^{+},\mu)}$. They then applied this estimate in an iteration scheme similar to the one we presented above to show the boundedness of $D_{d}u$ and $\mathcal{U}/x_{d}$. Due to the use of Moser’s iteration, their argument cannot be adapted to the case of systems with $n\geq 2$.
For $r\in(0,1]$ and $Q\subset\mathbb{R}^{d+1}$, we define the $(r/2,r)$-Hölder semi-norm of a function $f$ in $Q$ by
$f]_{C^{r/2,r}(Q)}=\sup_{(t,x),(s,y)\in Q,(t,x)\neq(s,y)}\frac{|f(t,x)-f(s,y)|}{|t-s|^{r/2}+|x-y|^{r}}.$
@@
where $N=N(d,n,\alpha,p,\kappa)>0$.
@@
where $N=N(d,n,\alpha,p,\kappa)>0$.
For $Q\subset\mathbb{R}^{d+1}_{+}$, we denote
$(f)_{Q}=\fint_{Q}fd\mu=\frac{1}{\mu(Q)}\int_{Q}fd\mu.$
@@
Let $p\in(1,2]$, $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r\in(0,\infty),\nu\in[32,\infty)$, and $u\in\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu)$ be a weak solution to (4.4)-(4.5) in $Q_{\nu r}^{+}(z_{0})$. Then
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}u-(D_{x^{\prime}}u)_{Q_{r}^{+}(z_{0})}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{U}-(\mathcal{U})_{Q_{r}^{+}(z_{0})}|d\mu$
$+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|u-(u)_{Q_{r}^{+}(z_{0})}|d\mu\leq N\nu^{-1}\left(\fint_{Q^{+}_{\nu r}(z_{0})}(|Du|^{p}+\lambda^{p/2}|u|^{p})d\mu\right)^{1/p},$
where $N=N(d,n,\alpha,p,\kappa)>0$.

### v2

@@ 4.1. Lipschitz estimates for homogeneous systems
$|D_{x^{\prime}}u|+\sqrt{\lambda}|u|\leq N_{d,\kappa}C_{2}.$
From the definition of $\mathcal{U}$ and the equation, we have $|\partial_{t}^{I}D_{x^{\prime}}^{J}\mathcal{U}|\leq N_{\kappa}|D\partial_{t}^{I}D_{x^{\prime}}^{J}u|$ and $|D_{d}\partial_{t}^{I}D_{x^{\prime}}^{J}\mathcal{U}|\leq N_{d,\kappa}(|\partial_{t}\partial_{t}^{I}D_{x^{\prime}}^{J}u|+|DD_{x^{\prime}}\partial_{t}^{I}D_{x^{\prime}}^{J}u|+\lambda|\partial_{t}^{I}D_{x^{\prime}}^{J}u|)$ for all $I,J\geq 0$. We then estimate the term $\mathcal{U}$ similarly as in the estimate of $u$ above, and get
$|\mathcal{U}|\leq N_{d,\kappa}C_{2}.$
Combining the above estimates and (4.6), we estimate $D_{d}u$ and arrive to (4.8) for $p=2$ and $\tau=1/2$.
By a covering argument, see, for example, [28, p. 184-186] or the proof of Lemma 4.5 below, estimates (4.8) hold for $p=2$ and general $\tau\in[1/2,1)$. Then by a standard iteration argument, see again [28, p. 184-186], estimates (4.8) hold for general $p\in(1,2]$, assuming that initially $u\in\mathcal{H}^{1}_{2}(Q_{r})$.
Assume now that $u\in\mathcal{H}^{1}_{p}(Q_{r})$, where $p\in(1,2)$, and let $\tau\in[1/2,1)$ be fixed. Let $u^{(\varepsilon)}$ be the mollification of $u$ in the $t,x^{\prime}$ variables. Let $r_{1}=\frac{1+\tau}{2}r$ and $r_{2}=\frac{3+\tau}{4}r$. As above,
$\|u^{(\varepsilon)}\|_{L_{\infty}(Q_{r_{1}})}^{p}\lesssim\int_{-r_{2}}^{r_{2}}(\|D_{d}u^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2,k}_{p}(Q^{\prime}_{r_{2}})}+\|u^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2,k}_{p}(Q^{\prime}_{r_{2}})})ds<\infty$
for a fixed large enough even $k=k(p,d)$. Then $\|D_{x^{\prime}}u^{(\varepsilon)}\|_{L_{\infty}(Q_{r_{1}})}<\infty$ as well. Next,
$\|\mathcal{U}^{(\varepsilon)}\|^{p}_{L_{\infty}(Q_{r_{1}})}$ $\lesssim\int_{-r_{2}}^{r_{2}}(\|D_{d}\mathcal{U}^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2,k}_{p}(Q^{\prime}_{r_{2}})}+\|\mathcal{U}^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2,k}_{p}(Q^{\prime}_{r_{2}})})ds$
$\lesssim\int_{-r_{2}}^{r_{2}}(\|D_{d}u^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2+1,k+2}_{p}(Q^{\prime}_{r_{2}})}+\|u^{(\varepsilon)}(\cdot,s)\|^{p}_{W^{k/2+1,k+2}_{p}(Q^{\prime}_{r_{2}})})ds<\infty,$
and therefore $\|D_{d}u^{(\varepsilon)}\|_{L_{\infty}(Q_{r_{1}})}<\infty$ as well. In particular, $u^{(\varepsilon)}\in\mathcal{H}^{1}_{2}(Q_{r_{1}})$ for all small enough $\varepsilon$. Then, (4.8) holds with $u^{(\varepsilon)}$ in place of $u$ in the left hand sides. Passing to the limit along a subsequence as $\varepsilon\to 0^{+}$, we obtain (4.8) for $u$. ∎
The following lemma is the key intermediate step for the proofs of our main results. Here we prove boundary reverse Hölder type $L_{\infty}$ estimates of solutions.
@@
Let $p\in(1,2]$, $r>0$, and $u\in\mathcal{H}^{1}_{p}(Q_{r}^{+},\mu;\mathbb{R}^{n})$ be a weak solution to (4.4)-(4.5) in $Q_{r}^{+}$. Then for any $\tau\in(0,1)$ it holds that
$\begin{gathered}\|u\|_{L_{\infty}(Q_{\tau r}^{+})}\leq N(1-\tau)^{-\frac{d+2+\alpha_{+}}{p}}\left(\fint_{Q_{r}^{+}}|u|^{p}d\mu\right)^{1/p},\\ \|Du\|_{L_{\infty}(Q_{\tau r}^{+})}+\sqrt{\lambda}\|u\|_{L_{\infty}(Q_{\tau r}^{+})}\leq N(1-\tau)^{-\frac{d+2+\alpha_{+}}{p}}\left(\fint_{Q_{r}^{+}}(|Du|^{p}+\lambda^{p/2}|u|^{p})d\mu\right)^{1/p},\\ \|\mathcal{U}/x_{d}\|_{L_{\infty}(Q_{\tau r}^{+})}\leq Nr^{-1}(1-\tau)^{-\frac{d+2+\alpha_{+}}{p}-1}\left(\fint_{Q_{r}^{+}}(|Du|^{p}+\lambda^{p/2}|u|^{p})d\mu\right)^{1/p},\end{gathered}$ (4.9)
where $N=N(d,\alpha,p,\kappa)>0$.
@@
It suffices to consider the case when $\tau\in[1/2,1)$.
We first consider the case $p=2$ and $\tau=1/2$. By scaling, it is sufficient to consider $r=1$. For compactness of notation, let
@@
Throughout the proof, the constants $N$ may depend on $d,\alpha,p$, and $\kappa$. Fix the smallest even integer $k$ with $k>(d+1)/2$. From the equation, we have $|D_{d}(x_{d}^{\alpha}\mathcal{U})|\leq Nx_{d}^{\alpha}(|u_{t}|+|DD_{x^{\prime}}u|+\lambda|u|)$. Then from $\lim_{x_{d}\to 0^{+}}x_{d}^{\alpha}\mathcal{U}=0$, Sobolev embedding, Hölder’s inequality, and Lemma 4.3, for $z\in Q_{1/2}^{+}$ it holds that
@@
$\leq NC_{1}+N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})\left(\int_{0}^{1/2}\|D_{d}u(\cdot,s)\|^{2}_{W_{2}^{k/2,k}(Q_{1/2}^{\prime})}s^{\alpha}ds\right)^{1/2}\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})C_{1}.$
In the above, estimate $|u(z^{\prime},1/2)|\leq NC_{1}$ follows from the interior $L_{\infty}$ estimate in Lemma 4.4, $\delta\in(0,1)$ is any constant when $\alpha\in\{1,3,5,\dots\}$, and $\delta=0$ otherwise. We introduce small $\delta>0$ to avoid the log-correction appearing in integrals $\int x_{d}^{-1}dx_{d}$. Applying estimate (4.11) to $u_{t},D_{x^{\prime}}u,D^{2}_{x^{\prime}}u,\lambda u$, and $\lambda^{1/2}u$, and using Lemma 4.3 give that on $Q_{1/2}^{+}$ we have
$\begin{gathered}|u_{t}|+|D_{x^{\prime}}u|+|D^{2}_{x^{\prime}}u|+\lambda|u|+\lambda^{1/2}|u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})C_{2}.\end{gathered}$ (4.12)
@@
$|D_{d}u|\leq N|\mathcal{U}|+N|D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})C_{2}.$ (4.13)
@@
$|D_{d}D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})C_{2}.$ (4.14)
@@
$\begin{gathered}|x_{d}^{\alpha}\mathcal{U}|\leq NC_{2}\int_{0}^{x_{d}}s^{\alpha}(1+s^{\frac{1-\alpha}{2}-\delta})ds\leq N(x_{d}^{1+\alpha}+x_{d}^{\frac{3+\alpha}{2}-\delta})C_{2},\end{gathered}$
@@
$|\mathcal{U}/x_{d}|\leq N(1+x_{d}^{\frac{1-\alpha}{2}-\delta})C_{2}.$
If $\alpha\in(-1,1)$, we are done with the case $p=2$. If $\alpha\geq 1$, we iterate the estimates above to improve the exponent $\frac{1-\alpha}{2}-\delta$ as follows. Plugging (4.13) into the first line of (4.11), we get
$|u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\delta})C_{1}.$
@@
$|u_{t}|+|D_{x^{\prime}}u|+|D^{2}_{x^{\prime}}u|+\lambda|u|+\lambda^{1/2}|u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\delta})C_{2}.$
@@
$|D_{d}u|+|D_{d}D_{x^{\prime}}u|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\delta})C_{2}.$
@@
$|\mathcal{U}/x_{d}|\leq N(1+x_{d}^{\frac{3-\alpha}{2}-\delta})C_{2}.$
@@
For general $\tau\in[1/2,1)$, we cover $Q_{\tau r}^{+}$ with boundary cylinders of the form $Q_{(1-\tau)r}^{+}(z_{0}^{\prime},0)$ with $z_{0}^{\prime}\in\overline{Q_{\tau r}^{\prime}}$, and interior cylinders $Q_{(1-\tau)r/4}(z_{0})$ with $z_{0}\in\overline{Q_{\tau r}^{+}}$ having $x_{0d}\geq(1-\tau)r/2$. On each boundary cylinder we apply the estimates proven above for $\tau=1/2$, and on each interior cylinder we use the estimates from Lemma 4.4. We then use inequality (4.7) for the measure $\mu$ to bound averages over the boundary and interior cylinders by the averages over $Q_{r}^{+}$. This completes the proof in the case $p=2$. Then by a standard iteration argument, see [28, p. 184-186], estimates (4.9) hold for general $p\in(1,2]$, assuming that initially $u\in\mathcal{H}^{1}_{2}(Q_{r}^{+},\mu)$.
Assume now that $u\in\mathcal{H}^{1}_{p}(Q_{r}^{+},\mu)$, where $p\in(1,2)$, and let $\tau\in[1/2,1)$ be fixed. Let $u^{(\varepsilon)}$ be the mollification of $u$ in the $t,x^{\prime}$ variables. Let $r_{1}=\frac{1+\tau}{2}r$ and $r_{2}=\frac{3+\tau}{4}r$. Note that the argument above can essentially be repeated by taking
$C_{1}=C_{2}=\left(\int_{0}^{r_{2}}\Bigl[\|Du^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q_{r_{2}}^{\prime})}^{p}+\|u^{(\varepsilon)}(\cdot,s)\|_{W^{k/2,k}_{p}(Q_{r_{2}}^{\prime})}^{p}\Bigr]s^{\alpha}ds\right)^{1/p}<\infty,$
where $k=k(p,d)$ is a fixed large enough even number. Indeed, by following the derivation of estimates above, the following sequence of estimates can be obtained for $z\in Q_{r_{1}}^{+}$. Let $\delta\in(0,1)$ be any constant if $\frac{1+\alpha}{p}\in\{1,2,3,\dots\}$ and $\delta=0$ otherwise. Then
@@
and so on. In finitely many steps, we get $|u^{(\varepsilon)}|+|Du^{(\varepsilon)}|+|\mathcal{U}^{(\varepsilon)}/x_{d}|\in L_{\infty}(Q_{r_{1}}^{+})$. In particular, $u^{(\varepsilon)}\in\mathcal{H}^{1}_{2}(Q_{r_{1}}^{+},\mu)$ for all small enough $\varepsilon$. Then, (4.9) holds with $u^{(\varepsilon)}$ in place of $u$ in the left hand sides. Passing to the limit along a subsequence as $\varepsilon\to 0^{+}$, we obtain (4.9) for $u$. ∎
@@
It should be noted that in the scalar case $n=1$ estimates (4.9) were proved in [18, pp. 1481–1487] with a somewhat different approach. There the authors first used Moser’s iteration to show the bound $\|u\|_{L_{\infty}(Q_{1/2}^{+})}\leq N\|u\|_{L_{2}(Q_{1}^{+},\mu)}$. They then applied this estimate in an iteration scheme similar to the one we presented above to show the boundedness of $D_{d}u$ and $\mathcal{U}/x_{d}$. Due to the use of Moser’s iteration, their argument cannot be adapted to the case of systems with $n\geq 2$.
For $r\in(0,1]$ and $Q\subset\mathbb{R}^{d+1}$, we define the $(r/2,r)$-Hölder semi-norm of a function $f:Q\to X$, where $X$ is a Banach space, by
$f]_{C^{r/2,r}(Q)}=\sup_{(t,x),(s,y)\in Q,(t,x)\neq(s,y)}\frac{|f(t,x)-f(s,y)|_{X}}{|t-s|^{r/2}+|x-y|^{r}}.$
@@
where $N=N(d,\alpha,p,\kappa)>0$.
@@
where $N=N(d,\alpha,p,\kappa)>0$.
@@
Let $p\in(1,2]$, $z_{0}\in\overline{\mathbb{R}^{d+1}_{+}},r\in(0,\infty),\nu\in[32,\infty)$, and $u\in\mathcal{H}^{1}_{p}(Q_{\nu r}^{+}(z_{0}),\mu;\mathbb{R}^{n})$ be a weak solution to (4.4)-(4.5) in $Q_{\nu r}^{+}(z_{0})$. Then
$\fint_{Q_{r}^{+}(z_{0})}|D_{x^{\prime}}u-(D_{x^{\prime}}u)_{Q_{r}^{+}(z_{0}),\mu}|d\mu+\fint_{Q_{r}^{+}(z_{0})}|\mathcal{U}-(\mathcal{U})_{Q_{r}^{+}(z_{0}),\mu}|d\mu$
$+\sqrt{\lambda}\fint_{Q_{r}^{+}(z_{0})}|u-(u)_{Q_{r}^{+}(z_{0}),\mu}|d\mu\leq N\nu^{-1}\left(\fint_{Q^{+}_{\nu r}(z_{0})}(|Du|^{p}+\lambda^{p/2}|u|^{p})d\mu\right)^{1/p},$
where $N=N(d,\alpha,p,\kappa)>0$.